OpenStudy (anonymous):

1. solve the equation using the domain of x = {1, 0, 2} 2x + 3y = 7 2. Give the range of x - 2y = 0 for the domain {-1, 0, 4} 3. The interest on a loan at an interest rate of 9% over a year is I = (0.09)P, where P is the principle (loan amount) and I is the interest rate. Use a domain in the table to find the interest of these amounts for one year. Write the range only in your solution. Principle: 10 100 1000 10,000 100,000 1,000,000 Interest: 4. What is the independent variable of this relation? Principle: 10 100 1,000 10,000 Interest: 2 20 200 2,000 Please help :)Thanks!

OpenStudy (anonymous):

1. first solve for any value of x. starting with 2x + 3y = 7 subtract 2x from both sides of the equation, then divide both sides of the equation by 3. you'll end up with: y = (-2x + 7)/3 now just evaluate for each specific value of x. when x=-1, y=(-2*-1 + 7)/3 = 9/3 = 3, so you have the point (-1, 3) when x=0, y=7/3, so (0, 7/3) when x=2, y=1, so (2, 1)

OpenStudy (anonymous):

2. x=-1 ---> y = -1/2 x= 0 ---> y = 0 x=4 --->y = 2 range = {-1/2 , 0 , 2}

OpenStudy (anonymous):

3. all you have to do is plug in P = 10 P = 100 P = 1000 etc into I = 0.09P to find the range

OpenStudy (anonymous):

tell me the answer in number 3. I have given you a solution

OpenStudy (anonymous):

@girlbyte so 3 would be 900%. But I'm still a bit stuck on #4

OpenStudy (anonymous):

3. answer is {0.9, 9, 90, 900, 9,000, 90,000}

OpenStudy (anonymous):

4. The principle is independent variable

OpenStudy (anonymous):

Okay, I totally get it now :3 Thanks so much! You've really helped, and I gave you a medal. @GirlByte

OpenStudy (anonymous):

@heirofkawaii thx. It would help me out alot if you fanned me too ;) haha